OpenStudy (anonymous):

Alg 1 Honors Help Please What is the exact value of the expression the square root 72. − the square root of 8. + the square root of 128.? Simplify if possible. 8the square root of 2. 12the square root of 2. 8the square root of 3. 12the square root of 3. What is the exact value of the expression the square root of 9 times y to the fourth power. + 12the square root of y to the sixth power. − 3the square root of y to the sixth power. − the square root of y to the fourth power.? Simplify if possible. 8y2 + 9y3 11y2 11y5 2y2 + 9y3

OpenStudy (anonymous):

@Hero @radar @jim_thompson5910 Can you guys help me please?

jimthompson5910 (jim_thompson5910):

what do you get when you simplify \[\large \sqrt{72}\]

OpenStudy (anonymous):

√9*8

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

all the possible ones for 72? or do you mean keep going with 8

jimthompson5910 (jim_thompson5910):

actually it would be better to write 72 as 36*2

jimthompson5910 (jim_thompson5910):

like this \[\large \sqrt{72} = \sqrt{36*2}\] \[\large \sqrt{72} = \sqrt{36}*\sqrt{2}\] \[\large \sqrt{72} = 6\sqrt{2}\]

OpenStudy (anonymous):

ohh! and for 8, \[2\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

good, you're getting the hang of this

jimthompson5910 (jim_thompson5910):

how about sqrt(128)

OpenStudy (anonymous):

For 128, \[2\sqrt{32}\]

jimthompson5910 (jim_thompson5910):

you can keep going

OpenStudy (anonymous):

\[1\sqrt{16}\]

jimthompson5910 (jim_thompson5910):

hmm not sure how you're getting that

jimthompson5910 (jim_thompson5910):

hint: 128 = 64*2

jimthompson5910 (jim_thompson5910):

the idea is to factor the number into 2 numbers where one of the numbers is the largest possible perfect square possible

OpenStudy (anonymous):

ohh! \[8\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

\[12\sqrt{2}\]

OpenStudy (anonymous):

is the final answer?

jimthompson5910 (jim_thompson5910):

so \[\large \sqrt{72} - \sqrt{8} + \sqrt{128}\] turns into \[\large 6\sqrt{2} - 2\sqrt{2} + 8\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

yep, then you combine like terms to get \[\large 6\sqrt{2} - 2\sqrt{2} + 8\sqrt{2} = 12\sqrt{2}\]

OpenStudy (anonymous):

Thank you so much! Can you help me with a few more problems please?

jimthompson5910 (jim_thompson5910):

sure, but just a few though

jimthompson5910 (jim_thompson5910):

what else do you need help with

OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 150r2 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r cubed square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

jimthompson5910 (jim_thompson5910):

here's a similar example

jimthompson5910 (jim_thompson5910):

let's say that each piece of paper covers an area of 10 square cm if the papers don't overlap and they cover the entire board that's 270 square cm then this means that you would need 270/10 = 27 sheets of paper

jimthompson5910 (jim_thompson5910):

put another way (Total Area of Board) = (# of sheets of paper)*(area of one sheet of paper) and you can solve/rearrange things to get (# of sheets of paper) = (Total Area of Board)/(area of one sheet of paper)

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Kind of

jimthompson5910 (jim_thompson5910):

what's not making sense

OpenStudy (anonymous):

yup i get it

jimthompson5910 (jim_thompson5910):

you sure?

OpenStudy (anonymous):

Yupp.

jimthompson5910 (jim_thompson5910):

ok so in our case, the total area is 150r^2 the area of each piece is sqrt( r^3 )

jimthompson5910 (jim_thompson5910):

if x = # of sheets of paper, then, (Total Area of Board) = (# of sheets of paper)*(area of one sheet of paper) ( 150r^2 ) = ( x )*( sqrt( r^3 ) ) 150r^2 = x *sqrt( r^3 ) now solve for x

OpenStudy (anonymous):

First, isolate the x. That makes the equation x=√(r^3) - 150 r^2

jimthompson5910 (jim_thompson5910):

no, but that's a good try

jimthompson5910 (jim_thompson5910):

the symbol * means "times"

jimthompson5910 (jim_thompson5910):

ex: 2 * 3 = 2 times 3 = 6

jimthompson5910 (jim_thompson5910):

so if you have something like 10 = x*2 then how do you isolate x?

OpenStudy (anonymous):

x=10/2

jimthompson5910 (jim_thompson5910):

good, you divide to undo multiplication

jimthompson5910 (jim_thompson5910):

so how can you apply that idea to 150r^2 = x *sqrt( r^3 ) so you can solve for x

OpenStudy (anonymous):

x=\[\frac{ 150r^2 }{ \sqrt{r^3} }\]

jimthompson5910 (jim_thompson5910):

very good, now rationalize the denominator

jimthompson5910 (jim_thompson5910):

and simplify

OpenStudy (anonymous):

r*r*r

jimthompson5910 (jim_thompson5910):

it might help to simplify sqrt(r^3) first

jimthompson5910 (jim_thompson5910):

doing so will give you \[\large \sqrt{r^3} = \sqrt{r^2*r}\] \[\large \sqrt{r^3} = \sqrt{r^2}*\sqrt{r}\] \[\large \sqrt{r^3} = r\sqrt{r}\]

jimthompson5910 (jim_thompson5910):

so \[\large x = \frac{150r^2}{ \sqrt{r^3} }\] becomes \[\large x = \frac{150r^2}{ r\sqrt{r} }\]

OpenStudy (anonymous):

yuppp

jimthompson5910 (jim_thompson5910):

then that turns into \[\large x = \frac{150r}{\sqrt{r} }\] I'll leave the last part to you

OpenStudy (anonymous):

x=150?

OpenStudy (anonymous):

I dont know how to divide with roots. :/

jimthompson5910 (jim_thompson5910):

how would you rationalize the denominator

OpenStudy (anonymous):

\[\frac{ \sqrt{r} }{ 150r}\]

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you need to multiply top and bottom by sqrt(r)

jimthompson5910 (jim_thompson5910):

then simplify

OpenStudy (anonymous):

\[\frac{ 150r }{ r }\times \frac{ \sqrt{r} }{ \sqrt{r}} \]

OpenStudy (anonymous):

oops forgot the √ for the bottom one

jimthompson5910 (jim_thompson5910):

that's ok, i know what you meant

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

I dont know how to multiply √r.

jimthompson5910 (jim_thompson5910):

what is √r times √r.

OpenStudy (anonymous):

√r^2

jimthompson5910 (jim_thompson5910):

which simplifies to what

jimthompson5910 (jim_thompson5910):

assume r > 0

OpenStudy (anonymous):

1√r?

jimthompson5910 (jim_thompson5910):

square roots undo squaring

OpenStudy (anonymous):

√r/1

jimthompson5910 (jim_thompson5910):

ex: 5 squared = 5^2 = 5*5 = 25 the square root of 25 = sqrt(25) = 5

OpenStudy (anonymous):

√r * √r

jimthompson5910 (jim_thompson5910):

so sqrt( x^2 ) = x assuming x > 0

jimthompson5910 (jim_thompson5910):

which means sqrt( r^2 ) = r where r > 0

jimthompson5910 (jim_thompson5910):

so we've gone from \[\large x = \frac{150r}{\sqrt{r} }\] to \[\large x = \frac{150r\sqrt{r}}{r }\] after multiplying top and bottom by sqrt(r)

jimthompson5910 (jim_thompson5910):

what's the last step

OpenStudy (anonymous):

x=150√r

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so you need \(\large 150\sqrt{r}\) sheets of paper

OpenStudy (anonymous):

YASYDYAYAYASYYy THANK YOU!!!:)

jimthompson5910 (jim_thompson5910):

lol glad I could be of help

OpenStudy (anonymous):

Aww Thank you so much for saving me.

jimthompson5910 (jim_thompson5910):

didn't do much really, but you're very welcome